Team:KULeuven/28 August 2009


Project progress

Progress of parts

[edit] Blue Light Receptor

  1. Miniprep of liquid cultures of pSB3K3 grown overnight.
Part concentration (ng/μl) 260/280 λ
pSB3K3 1 36,4 2,23
pSB3K3 2 33,0 2,46
  1. Gel electrophoresis of pSB3K3 miniprep and of restriction digest from 27/08 (BBa_J23101 and BBa_E0240. There was an unexpected signal at 800bp at the (BBa_J23101 lanes. The conclusion was that this part can not be used and we have to start all over from the basic DNA to make this ligation C.
  2. We checked our cultures that were radiated with blue light on 27/08.
    • The liquid cultures were put in the FACs machine. There was some fluorescence which is probably due to GFP. However, it can not be excluded that this is due to leakage from our promotor. So, a new set up needs to be made where decent controls are included.
    • The LB plates were put under a special GFP lamp. Just like the liquid cultures, there was signal but the same considerations need to be made.
  3. A new plan was designed in order to make a permanent glycerol stock of LigA and to make a new set up for GFP measurements, this time including controls.
    • LigA and BBa_E0240 were electroporated, plated and put overnight in the 37°C incubator.
    • BBa_E0240 was also plated starting from a glycerol stock and grown overnight at 37°.

The following was done over the weekend:

  • Saturday:
    • LigA grew well overnight. 4 colonies were selected for further testing and plated on new plates
    • The plates with BBa_E0240 overgrew, so a single colony was picked and replated
    • The plates containing electroporated ligC from 26/08 had two colonies after all. These were ented in liquid cultures and grown overnight, so it can be checked later if these colonies do in fact contain ligC.
  • Sunday:
    • The 4 colonies that were selected on Saturday grew well and were ented in liquid cultures.
    • Some colonies from BBa_E0240 were also ented in liquid culture.

[edit] Vanillin Production

  • We had to redo yesterday's test with the same volumes, allowing the gel to run longer
    • EF ligation was too short
    • Restriction digest for new ech and fcs parts (and also sam5 and sam8)
    • Put overnight in 16°C

[edit] Vanillin Receptor

  • new TOPO-cloning was done but with a change in protocol. Ligationtime of the fragement in the vector was changed from 5 minutes to overnight.
  • mutagenesis of rpoA on 202 was done but this showed no result on gel.
  • A new PCR was done to have more virA

[edit] Key/Lock/Anti-Key